A rifle that shoots bullets at 460 m/s is to be aimed at a target 45.7 m away. If the center of the target is level with the rifle, how high above the target must the rifle barrel be pointed so that the bullet hits dead center
1
Expert's answer
2020-04-17T10:47:32-0400

We have:


"\tan{\theta}=\frac{l}{R}"

The range is

"R=\frac{v^2\sin{2\theta}}{g}"

"45.7=\frac{460^2\sin{2\theta}}{9.8}"

"\theta=0.0606\degree"

Thus,

"\tan{0.0606\degree}=\frac{l}{45.7}"

"l=0.0483\ m"